OpenStudy (anonymous):

An object moves along the x axis according to the equation x(t) = (2.90t2 - 2.00t + 3.00) m. (a) Determine the average speed between t = 1.80 s and t = 3.00 s. m/s (b) Determine the instantaneous speed at t = 1.80 s and at t = 3.00 s. m/s (t = 1.80) m/s (t = 3.00) (c) Determine the average acceleration between t = 1.80 s and t = 3.00 s. m/s2 (d) Determine the instantaneous acceleration at t = 1.80 s and t = 3.00 s. m/s2 (t = 1.80) m/s2 (t = 3.00)

OpenStudy (anonymous):

What´s is the entire question?

OpenStudy (anonymous):

An object moves along the x axis according to the equation x(t) = (2.90t2 - 2.00t + 3.00) m. (a) Determine the average speed between t = 1.80 s and t = 3.00 s. m/s (b) Determine the instantaneous speed at t = 1.80 s and at t = 3.00 s. m/s (t = 1.80) m/s (t = 3.00) (c) Determine the average acceleration between t = 1.80 s and t = 3.00 s. m/s2 (d) Determine the instantaneous acceleration at t = 1.80 s and t = 3.00 s. m/s2 (t = 1.80) m/s2 (t = 3.00)

OpenStudy (anonymous):

what must i do with the given time values?

OpenStudy (anonymous):

You must derivate.

OpenStudy (anonymous):

The v(t) is simply the derrivative? How did you get m/s units of the derrivative?

OpenStudy (anonymous):

also, so let me see if im understanding_ the first derrivative gives me the velocity and the second derrivative gives me the acceleration?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Here it´s the letter A.

OpenStudy (anonymous):

Here it´s the letter B.

OpenStudy (anonymous):

Just plug the values t=1.80s and t=2.00 s in the equation and after use the formulae V

OpenStudy (anonymous):

More help on the way